[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Sep 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Sep 30 09:00:27 UTC 2017


HSU Shift Log for Fri Sep 29 2017 15:57:20 GMT-0400 (EDT)

Shift: H

Forecaster: L Avila

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Bonefish still slow. Doug took a look at it, but it still  
took several minutes to load satellite loops this afternoon.

AWIPS-2/NCP issues: None

Other Techincal issues: None

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170930/0e08f57a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list